OpenStudy (anonymous):

Determine whether the following set of points lies on a straight line (1,2,-1), (2,4,-3), (4,8,-7)

OpenStudy (anonymous):

Direction cosines? or maybe you can use the unit vector.

OpenStudy (anonymous):

Not sure.. Direction vector or displacement vector? How do you do this in steps?

OpenStudy (callisto):

I'm not sure if it is correct For AB, it is i +2j -2k if you can't see it clearly

OpenStudy (anonymous):

I thought that's how to do it. Thank you very much, it's correct.

OpenStudy (callisto):

You're welcome

OpenStudy (anonymous):

Find the slope between each set: AB, BC, and AC. If they are all equal, they are collinear. Slope = (y1 - y2) / (x1 - x2) AB = (4 - 1) / (3 - 1) = 3 / 2 BC = (10 - 4) / (7 - 3) = 6 / 4 = 3 / 2 AC = (10 - 1) / (7 - 1) = 9 / 6 = 3 / 2